2[4 – 2(3 – x)] – 1 = 4[2(4x – 3) + 7] – 25 I need help solving this. I have to once again complete 4 steps and I was sent the correct answer, but once again the steps did not match. Any help would be great.Step 1-“Simplify each side separately. step2-Isolate the variable term on one side,step 3“Isolate the variable (x= a number)and step 4-“Check. Substitute the proposed solution . Can someone please explain this to me ? I don't understand how to show the problem step by step like that.''.

Answer by stanbon(75887)

2[4 – 2(3 – x)] – 1 = 4[2(4x – 3) + 7] – 25
Simplify the inner parentheses:
2[4 – 6+2x] – 1 = 4[8x-6 + 7] – 25
Combine like terms in the parentheses:
2[-2+2x] -1 = 4[8x+1] - 25
Simplify the parentheses again:
-4+4x-1 = 32x+4-25
Combine like terms on each side:
4x-3 = 32x-21
Subtract 4x from each side to get:
-3 = 28x-21
Add 21 to each side to get
28x = 18
Divide each side by 28 to get
x = 18/28
Reduce the fraction:
x = 9/14
================
Cheers,
Stan H.